# Decomposing The Dark Matter of Sparse Autoencoders

### v1

@@ Abstract
Sparse autoencoders (SAEs) are a promising technique for decomposing language model activations into interpretable linear features. However, current SAEs fall short of completely explaining model performance, resulting in “dark matter”: unexplained variance in activations. This work investigates dark matter as an object of study in its own right. Surprisingly, we find that much of SAE dark matter—about half of the error vector itself and $>90\%$ of its norm—can be linearly predicted from the initial activation vector. Additionally, we find that the scaling behavior of SAE error norms at a per token level is remarkably predictable: larger SAEs mostly struggle to reconstruct the same contexts as smaller SAEs. We build on the linear representation hypothesis to propose models of activations that might lead to these observations, including postulating a new type of “introduced error”; these insights imply that the part of the SAE error vector that cannot be linearly predicted (“nonlinear” error) might be fundamentally different from the linearly predictable component. To validate this hypothesis, we empirically analyze nonlinear SAE error and show that 1) it contains fewer not yet learned features, 2) SAEs trained on it are quantitatively worse, 3) it helps predict SAE per-token scaling behavior, and 4) it is responsible for a proportional amount of the downstream increase in cross entropy loss when SAE activations are inserted into the model. Finally, we examine two methods to reduce nonlinear SAE error at a fixed sparsity: inference time gradient pursuit, which leads to a very slight decrease in nonlinear error, and linear transformations from earlier layer SAE outputs, which leads to a larger reduction.
@@ 1 Introduction
The ultimate goal for ambitious mechanistic interpretability is to understand neural networks from the bottom up by breaking them down into programs (“circuits”) and the variables (“features”) that those programs operate on (Olah, 2023). One recent successful technique for finding features in language models has been sparse autoencoders (SAEs), which learn a dictionary of one-dimensional representations that can be sparsely combined to reconstruct model hidden activations (Cunningham et al., 2023; Bricken et al., 2023). However, as observed by Gao et al. (2024), the scaling behavior of SAE width (number of latents) vs. reconstruction mean squared error (MSE) is best fit by a power law with a constant error term. Gao et al. (2024) speculate that this component of SAE error below the asymptote might best be explained by model activations having components with denser structure than simple SAE features (e.g. Gaussian noise). This is a concern for the ambitious agenda because it implies that there are components of model hidden states that are harder for SAEs to learn and which might not be eliminated by simple scaling of SAEs.
@@
    In Section 5, we propose models of SAE error to explain these observations, including postulating a new type of “introduced error” that is caused by the SAE architecture and sparsity constraint.
@@
    In Section 6, we investigate the linearly predictable and nonlinearly predictable components of SAE error. We find that although the nonlinear component affects downstream cross entropy loss in proportion to its norm, it is qualitatively different from linear error: as compared to linear error, nonlinear error is harder to learn SAEs for, consists of a smaller proportion of absent linear features, and helps with predicting the norm of larger SAE errors from smaller SAEs.
  4. 4.
    In Section 7, we show that inference time optimization increases the fraction of variance explained by SAEs, but only slightly decreases nonlinear error. Additionally, we show that we can use SAEs trained on previous components to decrease nonlinear error and total SAE error.
Figure 1: A breakdown of SAE dark matter. See Section 4 for how we break down the overall fraction of unexplained variance into absent features, linear error, and nonlinear error. See Section 7.1 for further separating encoder error from nonlinear error.
@@ 2 Related Work
Language Model Representation Structure: The linear representation hypothesis (LRH) (Park et al., 2023; Elhage et al., 2022) claims that language model hidden states can be decomposed into a sparse sum of linear feature directions. The LRH has seen recent empirical support with sparse autoencoders, which have succeeded in decomposing much of the variance of language model hidden states into such a sparse sum, as well as a long line of work that has used probing and dimensionality reduction to find causal linear representations for specific concepts (Alain, 2016; Nanda et al., 2023; Marks et al., 2024; Gurnee, 2024). On the other hand, some recent work has questioned whether the linear representation hypothesis is true: Engels et al. (2024) find multidimensional circular representations in Mistral (Jiang et al., 2023) and Llama (AI@Meta, 2024), and Csordás et al. (2024) examine synthetic recurrent neural networks and find “onion-like” non-linear features not contained in a linear subspace. This has inspired recent discussion about what a true model of activation space might be: Mendel (2024) argues that the linear representation hypothesis ignores the growing body of results showing the multi-dimensional structure of SAE latents, and Smith (2024b) argues that we only have evidence for a “weak” form of the superposition hypothesis holding that only some features are linearly represented.
@@ 3 Notation
In this paper, we consider neural network activations ${\mathbf{x}}\in\mathbb{R}^{d}$ and sparse autoencoders $\texttt{Sae}\in\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ which seek to minimize $\left\lVert{\mathbf{x}}-\texttt{Sae}({\mathbf{x}})\right\rVert_{2}$ while using a small number of active latents. We are agnostic to the architecture or training procedure of the sparse autoencoder; see (Bricken et al., 2023; Cunningham et al., 2023; Gao et al., 2024; Templeton et al., 2024) for such details. We also define $\texttt{SaeError}({\mathbf{x}})$ such that ${\mathbf{x}}=\texttt{Sae}({\mathbf{x}})+\texttt{SaeError}({\mathbf{x}})$, or rearranged
$\texttt{SaeError}({\mathbf{x}})\coloneq{\mathbf{x}}-\texttt{Sae}({\mathbf{x}}).$ (1)
@@ 4 Predicting SAE Error
Experiment Details: [^1] Code at https://github.com/JoshEngels/SAE-Dark-Matter Unless noted otherwise, we set ${\mathbf{x}}$ equal to layer $20$ Gemma 2 9B (Team et al., 2024) activations. We use $300$ contexts of $1024$ tokens from the uncopywrited subset of the Pile (Gao et al., 2020) and then filter to only activations of tokens after position $200$ in each context, as Lieberum et al. (2024) find that earlier tokens are easier for sparse autoencoders to reconstruct, and we wish to ignore the effect of token position on our results. This results in a dataset of about 247k activations. We use the suite of Gemma Scope (Lieberum et al., 2024) sparse autoencoders. For linear regressions, we use a random subset of size 150k as training examples (since Gemma 2 9B has a hidden dimension of $3584$, this prevents overfitting) and report the $R^{2}$ on the other 97k activations. For transforms to a multi-dimensional output, we report the average $R^{2}$ across dimensions. We include bias terms in our linear regressions but omit them from equations for simplicity.
Figure 2: $R^{2}$ of linear regressions from activations to SAE error norms. Errors norms are linearly predictable with high accuracy at all widths and $L_{0}$.
Figure 3: $R^{2}$ of linear transformations from activations to SAE error vectors. Errors vector prediction accuracy decreases with increasing width and $L_{0}$.
Predicting SAE Error Norm: For our first set of experiments, we find the optimal linear probe ${\bm{a}}^{*}$ from ${\mathbf{x}}$ to $\lVert\texttt{SaeError}({\mathbf{x}})\rVert^{2}_{2}$. Formally (with a slight abuse of notation, since ${\mathbf{x}}$ is a random variable and not a dataset), we solve for
@@
The $R^{2}$ of these probes are all extremely high: on layer $20$, for all Gemma Scope combinations of SAE width and $L_{0}$, between $86\%$ and $95\%$ of the variance in SAE error norm is explained by the optimal linear probe. We plot these results as a contour plot in Fig. 3. Overall, sparser and wider SAEs have less predictable error norms. In Appendix A, we show a plot of the $R^{2}$ of the prediction across layers, and find that except for the first few layers, activation probes have a much higher $R^{2}$ than probes using tokens, SAE $L_{0}$, model loss, or activation norm.
Predicting SAE Error Vectors: We next examine the $R^{2}$ of the optimal linear transform ${\bm{b}}^{*}$ from ${\mathbf{x}}$ to $\texttt{SaeError}({\mathbf{x}})$:
@@
As we show in Fig. 3, the $R^{2}$ of these transforms for layer $20$ range between $30\%$ and $72\%$; this is less than the $R^{2}$ for our norm prediction experiments, but still much higher than we might expect. Intuitively, this result implies that there must be large linear subspaces that the SAE is mostly failing to learn. Like SAE error norm predictions, there is a clear pattern across SAE $L_{0}$ and width: $R^{2}$ decreases with increasing SAE width and $L_{0}$. Interestingly, this pattern is not the same as it was above for SAE error norm: the $R^{2}$ of error norm predictions increases with SAE $L_{0}$, while it decreases for error vector predictions. One concern might be that ${\bm{b}}^{*}$ is mostly reversing feature shrinkage, in Section A.1, we show that this is not the case.
Nonlinear FVU: Another related metric we are interested in is the total amount of the original activation ${\mathbf{x}}$ we fail to “explain” using both the SAE reconstruction $\texttt{Sae}({\mathbf{x}})$ and a linear projection of ${\mathbf{x}}$. That is, assuming we have found ${\bm{b}}^{*}$ as in Eq. 3, we are interested in the fraction of variance unexplained (FVU) by the sum of $\texttt{Sae}({\mathbf{x}})$ and ${\bm{b}}^{*}\cdot{\mathbf{x}}$:
@@
We label this quantity $\texttt{FVU}_{\texttt{nonlinear}}$ because it is intuitively the amount of the SAE’s unexplained variance that is not a linear projection of the input. Interestingly, we find that at a fixed $L_{0}$, $\texttt{FVU}_{\texttt{nonlinear}}$ is approximately constant (see Fig. 5). That is, even though we can linearly predict a smaller portion of the error vector in larger SAEs, this effect is counteracted almost exactly by the fact that the SAE error vector itself is getting smaller. In contrast, $\texttt{FVU}_{\texttt{nonlinear}}$ decreases as SAE $L_{0}$ increases.
We can use the hypothesis that $\texttt{FVU}_{\texttt{nonlinear}}$ is a constant at a fixed sparsity to plot the breakdown of FVU for varying SAE width and $L_{0}\approx 60$ in Fig. 1. We plot a horizontal fit for $\texttt{FVU}_{\texttt{nonlinear}}$ and a power law fit with a constant for Gemma SAE reconstructions; the power law fit asymptotes above the horizontal fit, which implies the presence of linear error even at very large SAE width. Note that we assume absent features (those which are not yet learned) are a component of the linear error; we provide justifications for this in Section 5.2.
Figure 4: $\texttt{FVU}_{\texttt{nonlinear}}$ versus SAE width and $L_{0}$. Larger SAE $L_{0}$s have a smaller $\texttt{FVU}_{\texttt{nonlinear}}$, but $\texttt{FVU}_{\texttt{nonlinear}}$ stays mostly constant with increasing SAE width.
@@
Predicting SAE Per-Token Error Norms: We now examine per-token SAE scaling behavior. Given two SAEs, $\texttt{SAE}_{1}$ and $\texttt{SAE}_{2}$, we are interested in how much of the variance in error norms in $\texttt{SAE}_{2}$ is predictable from error norms in $\texttt{SAE}_{1}$. That is, we want to find ${\bm{c}}^{*}$ such that
${\bm{c}}^{*}\coloneq\operatorname*{arg\,min}_{{\bm{c}}\in\mathbb{R}}\left\lVert{\bm{c}}\cdot\texttt{SaeError}_{1}({\mathbf{x}})-\texttt{SaeError}_{2}({\mathbf{x}})\right\rVert_{2}$ (5)
Note that in practice, although ${\bm{a}}^{*}$ also can predict the norm of SAE error, it requires training the target SAE to learn a probe. Here, on the other hand, although we formulate finding ${\bm{c}}^{*}$ as an optimization problem that requires a larger SAE, in practice we do not need to actually train the larger SAE to get interesting insights: since ${\bm{c}}^{*}$ has just one component, it simply measures how well small SAE error can be multiplied by a scalar to predict large SAE error. If the $R^{2}$ is high, we know that on tokens that small SAEs perform poorly on, larger SAEs will as well. In Fig. 5, we plot the $R^{2}$ of ${\bm{c}}^{*}$ probes on all pairs of layer $20$ SAEs with $L_{0}\approx 60$ (restricting to pairs where $\texttt{SAE}_{2}$ is larger than $\texttt{SAE}_{2}$), and find that indeed, per token SAE errors are highly predictable. Additionally, we show concretely what these correlated SAE error norms looks like on a set of $100$ tokens from the Pile in Fig. 6.

### v2

@@ Abstract
Sparse autoencoders (SAEs) are a promising technique for decomposing language model activations into interpretable linear features. However, current SAEs fall short of completely explaining model performance, resulting in “dark matter”: unexplained variance in activations. This work investigates dark matter as an object of study in its own right. Surprisingly, we find that much of SAE dark matter—about half of the error vector itself and $>90\%$ of its norm—can be linearly predicted from the initial activation vector. Additionally, we find that the scaling behavior of SAE error norms at a per token level is remarkably predictable: larger SAEs mostly struggle to reconstruct the same contexts as smaller SAEs. We build on the linear representation hypothesis to propose models of activations that might lead to these observations. These insights imply that the part of the SAE error vector that cannot be linearly predicted (“nonlinear” error) might be fundamentally different from the linearly predictable component. To validate this hypothesis, we empirically analyze nonlinear SAE error and show that 1) it contains fewer not yet learned features, 2) SAEs trained on it are quantitatively worse, and 3) it is responsible for a proportional amount of the downstream increase in cross entropy loss when SAE activations are inserted into the model. Finally, we examine two methods to reduce nonlinear SAE error: inference time gradient pursuit, which leads to a very slight decrease in nonlinear error, and linear transformations from earlier layer SAE outputs, which leads to a larger reduction.
@@ 1 Introduction
The ultimate goal for ambitious mechanistic interpretability is to understand neural networks from the bottom up by breaking them down into programs (“circuits") and the variables (“features”) that those programs operate on (Olah, 2023). One recent successful technique for finding features in language models has been sparse autoencoders (SAEs), which learn a dictionary of one-dimensional representations that can be sparsely combined to reconstruct model hidden activations (Cunningham et al., 2023; Bricken et al., 2023). However, as observed by Gao et al. (2024), the scaling behavior of SAE width (number of latents) vs. reconstruction mean squared error (MSE) is best fit by a power law with a constant error term. Gao et al. (2024) speculate that this component of SAE error below the asymptote might best be explained by model activations having components with denser structure than simple SAE features (e.g. Gaussian noise). This is a concern for the ambitious agenda because it implies that there are components of model hidden states that are harder for SAEs to learn and which might not be eliminated by simple scaling of SAEs.
@@
    In Section 5, we investigate the linearly predictable and nonlinearly predictable components of SAE error. We find that although the nonlinear component affects downstream cross entropy loss in proportion to its norm, it is qualitatively different from linear error: as compared to linear error, nonlinear error is harder to learn SAEs for and has a norm that is harder to linearly predict from activations, suggesting that it consists of a smaller proportion of not-yet-learned linear features.
@@
    In Section 6, we investigate methods for reducing nonlinear error. We show that inference time optimization increases the fraction of variance explained by SAEs, but only slightly decreases nonlinear error. Additionally, we show that we can use SAEs trained on previous components to decrease nonlinear error and total SAE error.
Figure 1: A breakdown of SAE dark matter for layer 20 Gemma 9B SAEs, with dotted lines assuming that observed trends continue for larger SAEs. See Section 4 for how we break down the overall fraction of unexplained variance into absent features, linear error, and nonlinear error. See Section 6.1 for further separating encoder error from nonlinear error.
@@ 2 Related Work
Language Model Representation Structure: The linear representation hypothesis (LRH) (Park et al., 2023; Elhage et al., 2022) claims that language model hidden states can be decomposed into a sparse sum of linear feature directions. The LRH has seen recent empirical support with sparse autoencoders (Makhzani & Frey, 2013; Bricken et al., 2023; Cunningham et al., 2023), which have succeeded in decomposing much of the variance of language model hidden states into such a sparse sum, as well as a long line of work that has used probing and dimensionality reduction to find causal linear representations for specific concepts (Alain, 2016; Nanda et al., 2023; Marks et al., 2024; Gurnee, 2024). On the other hand, some recent work has questioned whether the linear representation hypothesis is true: Engels et al. (2024) find multidimensional circular representations in Mistral (Jiang et al., 2023) and Llama (AI@Meta, 2024), and Csordás et al. (2024) examine synthetic recurrent neural networks and find “onion-like” non-linear features not contained in a linear subspace. This has inspired recent discussion about what a true model of activation space might be: Mendel (2024) argues that the linear representation hypothesis ignores the growing body of results showing the multi-dimensional structure of SAE latents, and Smith (2024b) argues that we only have evidence for a “weak” form of the superposition hypothesis holding that only some features are linearly represented (such a hypothesis has also been studied in the sparse coding literature (Tasissa et al., 2020)).
@@ 3 Notation
We consider neural network activations ${\mathbf{x}}\in\mathbb{R}^{d}$ and sparse autoencoders $\texttt{Sae}\in\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Sparse autoencoders map inputs to a latent space $\mathbb{R}^{m}$ with $m\gg d$ and then back into $\mathbb{R}^{d}$, while also requiring that only a sparse set of the $m$ latent dimensions are nonzero. SAEs have the general architecture
$\texttt{hidden}({\mathbf{x}})$ :-$\coloneq\sigma({\bm{W}}_{enc}\cdot({\mathbf{x}}-{\bm{b}}_{dec})+{\bm{b}}_{enc}):$ (1)
$\texttt{Sae}({\mathbf{x}})$ :-$\coloneq{\bm{W}}_{dec}\cdot\texttt{hidden}(x)+{\bm{b}}_{dec}:$ (2)
where $\sigma$ is an architecture dependent activation function, and seek to minimize
$\mathcal{L}_{\texttt{Sae}({\mathbf{x}})}=\left\lVert{\mathbf{x}}-\texttt{Sae}({\mathbf{x}})\right\rVert_{2}^{2}+S(\texttt{hidden}({\mathbf{x}}))$ (3)
where $S$ is an architecture dependent sparsity function. For convenience, we define $\texttt{SaeError}({\mathbf{x}})$ as the reconstruction error of the SAE and $L_{0}$ as the average number of nonzeros in $\texttt{hidden}({\mathbf{x}})$:
$\texttt{SaeError}({\mathbf{x}})$ :-$\coloneq{\mathbf{x}}-\texttt{Sae}({\mathbf{x}}):$ (4)
$L_{0}$ :-$\coloneq E_{\mathbf{x}}(\left\lVert\texttt{hidden}({\mathbf{x}})\right\rVert_{0}):$ (5)
In this work, we study the relationship between ${\mathbf{x}}$ and $\texttt{SaeError}({\mathbf{x}})$ and are agnostic to the specifics of SAE architecture. Thus, we examine both TopK SAEs (Gao et al., 2024) and JumpRelu SAEs (Rajamanoharan et al., 2024). The TopK SAE is defined as
$\sigma_{TopK}({\mathbf{z}})$ $=\text{set all but the $k$ largest dimensions of ${\mathbf{z}}$ to zero}$ (6)
$S_{TopK}(\texttt{hidden}({\mathbf{x}}))$ $=0$ (7)
Note that $L_{0}=k$ for a TopK SAE. The JumpRelu SAE is defined as
$\sigma_{JumpRelu}({\mathbf{z}})_{i}$ =$=\begin{cases}{\mathbf{z}}_{i}&\text{if ${\mathbf{z}}_{i}-\theta_{i}>0$}\\ 0&\text{otherwise}\end{cases}$ (8)
$S_{JumpRelu}(\texttt{hidden}({\mathbf{x}}))$ $=\lambda\left\lVert\texttt{hidden}({\mathbf{x}})\right\rVert_{0}$ (9)
where $\theta$ is a learned bias vector and $\lambda$ is a sparsity parameter.
@@ 4 Predicting SAE Error
In this section, we evaluate the extent to which SAE errors $\texttt{SaeError}({\mathbf{x}})$ can be predicted from input model activations ${\mathbf{x}}$.
We run experiments  [^1] Code at https://anonymous.4open.science/r/SAE-Dark-Matter-1163 on Gemma 2 2B and 9B (Team et al., 2024) and Llama 3.1 8B (AI@Meta, 2024). We use the suite of Gemma Scope (Lieberum et al., 2024) sparse autoencoders for Gemma 2 2B experiments and Llama Scope (He et al., 2024) for Llama 3.1 8B. For experiments where we analyze the effect of $L_{0}$ and SAE width, we focus on layer 12 of Gemma 2 2B and layer 20 of Gemma 9 9B in the main text as these layers have the most SAEs in Gemma Scope, and we include additional layers in the appendix (we do not run on Llama Scope for these experiments, as Llama Scope has just one SAE $L_{0}$ and two SAE widths per layer). For experiments where we analyze the effect of layer, we show the set of SAEs with $L_{0}$ closest to a target $L_{0}$ across all layers for both Gemma Scope and Llama Scope suites of SAEs.
We use $300$ contexts of $1024$ tokens from the uncopywrited subset of the Pile (Gao et al., 2020) and then filter to only activations of tokens after position $200$ in each context, as Lieberum et al. (2024) find that earlier tokens are easier for sparse autoencoders to reconstruct, and we wish to ignore the effect of token position on our results. This results in a dataset of about 247k activations. For linear regressions, we use a random subset of size 150k as training examples (since all models have a dimension of less than $5000$, this prevents overfitting) and report the $R^{2}$ on the other 97k activations. For linear transformations to a multi-dimensional output, we report the average $R^{2}$ across dimensions. We include bias terms in our linear regressions but omit them from equations for simplicity.
(a) Linear prediction results for layer 12 Gemma 2 2B SAEs from Gemma Scope. $\texttt{FVU}_{\texttt{nonlinear}}$ is roughly constant for a fixed $L_{0}$.
(b) Linear prediction results for layer 20 Gemma 2 9B SAEs from Gemma Scope. $\texttt{FVU}_{\texttt{nonlinear}}$ is roughly constant for a fixed $L_{0}$.
Figure 2: Results of linearly predicting SAE error norm and SAE error from model activations on Gemma 2 2B layer 12 (top) and Gemma 2 9B layer 20 (bottom). The right plots show the $R^{2}$ of predicting SAE error norms (see Eq. 10, the middle plots show the $R^{2}$ of predicting SAE error vectors (see Eq. 11, and the right plots show $1-R^{2}$ of predicting model activations given the SAE reconstruction and the SAE error vector prediction. We note that these 2D heatmaps are somewhat sparse and only the black dots represent actual SAEs. This is because we use Gemma Scope SAEs, which are trained only on some $L_{0}$s and SAE widths. We do a linear interpolation between SAEs to predict $R^{2}$ between hyperparameters.
4.1 Predicting SAE Error Norm
For our first set of experiments, we find the optimal linear probe ${\bm{a}}^{*}$ from ${\mathbf{x}}$ to $\lVert\texttt{SaeError}({\mathbf{x}})\rVert^{2}_{2}$. Formally (with a slight abuse of notation, since ${\mathbf{x}}$ is a random variable and not a dataset), we solve for
@@ 4.1 Predicting SAE Error Norm
The $R^{2}$ of these probes are all extremely high: across all combinations of SAE width, $L_{0}$, layer, and model, between $70\%$ and $95\%$ of the variance in SAE error norm is explained by the optimal linear probe.
Results across SAE width and $L_{0}$: We plot the $R^{2}$ of probes from Eq. 10 across SAE width and SAE $L_{0}$ for layer 9 of Gemma 2 2B and layer 20 of Gemma 2 9B as a contour plot (on the left) in Fig. 2. We also include plots for additional layers in Appendix A. Overall, sparser and wider SAEs have less predictable error norms, although interestingly for some layers, extremely dense SAEs have less predictable error norm as well.
Results across layers and models: In Fig. 4 we plot the $R^{2}$ of probes from Eq. 10 for the SAEs with $L_{0}$ closest to $50$ on Gemma 2 2B and 9B and Llama Scope SAEs (which all have $k=L_{0}=50$). We find that the $R^{2}$ is low for the first few layers, increases rapidly to a value of about $0.8$ to $0.95$, remains at this level with a slight dip for mid-late layers, and then drops off steeply in the last few layers. In Appendix A, we compare the $R^{2}$ of this approach across layers versus other baselines like token identity; we find that using activations does much better, and thus it is not “easy” to predict error norm.
4.2 Predicting SAE Error Vectors
We next examine the $R^{2}$ of the optimal linear transform ${\bm{b}}^{*}$ from ${\mathbf{x}}$ to $\texttt{SaeError}({\mathbf{x}})$:
@@ 4.2 Predicting SAE Error Vectors
As we show in Fig. 2 (middle column) and in Appendix A for additional layers, the $R^{2}$ of these transforms varies widely, between $15\%$ and $72\%$; this is less than the $R^{2}$ for our norm prediction experiments, but still much higher than we might expect. Intuitively, this result implies that there are large linear subspaces that the SAE is mostly failing to learn. There is also a clear pattern across SAE $L_{0}$ and width: $R^{2}$ decreases with increasing SAE width and $L_{0}$. Interestingly, this pattern is not the same as it was above for SAE error norm: the $R^{2}$ of error norm predictions increases with SAE $L_{0}$, while it decreases for error vector predictions. One concern might be that ${\bm{b}}^{*}$ is mostly reversing feature shrinkage; in Section A.1, we show that this is not the case.
4.3 Nonlinear FVU and Breaking Down SAE Dark Matter
Another related metric we are interested in is the total amount of the original activation ${\mathbf{x}}$ we fail to explain using both $\texttt{Sae}({\mathbf{x}})$ and a linear projection of ${\mathbf{x}}$. That is, assuming we have found ${\bm{b}}^{*}$ as in Eq. 11, we are interested in the fraction of variance unexplained (FVU) by the sum of $\texttt{Sae}({\mathbf{x}})$ and ${\bm{b}}^{*}\cdot{\mathbf{x}}$:
@@ 4.3 Nonlinear FVU and Breaking Down SAE Dark Matter
We label this quantity $\texttt{FVU}_{\texttt{nonlinear}}$ because it is intuitively the amount of the SAE’s unexplained variance that is not a linear projection of the input. Interestingly, we find that for middle layers (layer 12 Gemma 2 2B and layer 20 Gemma 2 9B) at a fixed $L_{0}$, $\texttt{FVU}_{\texttt{nonlinear}}$ is approximately constant (see Fig. 2), while for other layers it decreases slightly as $L_{0}$ increases (see Appendix A). That is, even though we can linearly predict a smaller portion of the error vector in larger SAEs, this effect is counteracted by the fact that the SAE error vector itself is getting smaller. In contrast, $\texttt{FVU}_{\texttt{nonlinear}}$ decreases as SAE $L_{0}$ increases.
As stated above, for middle layers we observe that $\texttt{FVU}_{\texttt{nonlinear}}$ is roughly constant at a fixed sparsity. Because we observe this across multiple orders of magnitude of SAE scale, we hypothesize that this will continue to hold as we scale the SAE. Using this assumption, we plot a horizontal fit for $\texttt{FVU}_{\texttt{nonlinear}}$ and a power law fit with a constant for Gemma SAE reconstructions as SAE width scales (choosing the SAEs closest to $L_{0}\approx 60$) in Fig. 1. The power law fit asymptotes above the horizontal fit, which implies the presence of linear error even at very large SAE width. The absent features component of the error in Fig. 1 comes from the following observation: if our hypothesis is correct that nonlinear error is roughly constant as the layer 20 SAEs scale, then the nonlinear features must reside in the linear error component, since this is the only component that decreases as the SAE scales and learns more features. We investigate this hypothesis further in Section B.2.
Figure 3: $R^{2}$ of SAE error norm predictions (see Eq. 10) for Gemma Scope SAEs of width $L_{0}\approx 50$ and Llama Scope SAEs of width $L_{0}=k=50$.
@@
4.4 Predicting SAE Per-Token Error Norms
We now examine per-token SAE scaling behavior. Given two SAEs, $\texttt{SAE}_{1}$ and $\texttt{SAE}_{2}$, we are interested in how much of the variance in error norms in $\texttt{SAE}_{2}$ is predictable from error norms in $\texttt{SAE}_{1}$. That is, we want to find ${\bm{c}}^{*}$ such that
${\bm{c}}^{*}\coloneq\operatorname*{arg\,min}_{{\bm{c}}\in\mathbb{R}}\left\lVert{\bm{c}}\cdot\left\lVert\texttt{SaeError}_{1}({\mathbf{x}})\right\rVert_{2}^{2}-\left\lVert\texttt{SaeError}_{2}({\mathbf{x}})\right\rVert_{2}^{2}\right\rVert_{2}$ (13)
Note that in practice, although ${\bm{a}}^{*}$ also can predict the norm of SAE error, it requires training the target SAE to learn a probe. Here, on the other hand, although we formulate finding ${\bm{c}}^{*}$ as an optimization problem that requires a larger SAE, in practice we do not need to actually train the larger SAE to get interesting insights: since ${\bm{c}}^{*}$ has just one component, it simply measures how well small SAE error can be multiplied by a scalar to predict large SAE error. If the $R^{2}$ is high, we know that on tokens that small SAEs perform poorly on, larger SAEs will as well. In Fig. 4, we plot the $R^{2}$ of ${\bm{c}}^{*}$ probes on all pairs of Gemma Scope 9B layer $20$ SAEs with $L_{0}\approx 60$ (restricting to pairs where $\texttt{SAE}_{2}$ is larger than $\texttt{SAE}_{2}$), and find that indeed, per token SAE errors are highly predictable. Additionally, we show concretely what these correlated SAE error norms looks like on a set of $100$ tokens from the Pile in Fig. 5.